OpenStudy (anonymous):

What is the Taylor Series for x^1/2 centered at a=16? I think this is something to do with a binomial series.

OpenStudy (anonymous):

i can grind it out, not sure if that is the best method

OpenStudy (anonymous):

Ill post what I got once the file sends over!

OpenStudy (kirbykirby):

A Taylor series for a function \(f(x)\) is expanded about a point \(x_0\) as follows \[\begin{align*} f(x) &= f(x_0) + f'(x_0)(x-x_0)+\frac{f''(x_0)}{2!}(x-x_0)^2 +\frac{f'''(x_0)}{3!}(x-x_0)^3+...\\ &=\sum_{n=0}^{\infty}\frac{f^{(n)}(x_0)}{n!}(x-x_0)^n \end{align*}\]

OpenStudy (anonymous):

I got the expansion that you said kirbykirby through a Binomial series, which i am not sure if it is right or not. I seem to be having some issues with the denominator and the factorials matching up, in particular the \[\frac{ -x^2 }{8 }\] part. The factorial or denominator does not seem to be making a pattern, which means I probably made a really stupid error somewhere, unless I did it totally wrong. In which case, does anyone have any suggestions?

OpenStudy (anonymous):

your terms should look like \(x-16\)

OpenStudy (anonymous):

think you are going to use this, right \[(1+x)^{\frac{1}{2}}=\sum_{n\ge 0}\binom{\frac{1}{2}} n x^n\]

OpenStudy (anonymous):

Yes, where the 1 on the left is substituted with a 16, correct?

OpenStudy (anonymous):

and you are going to have to change your variable

OpenStudy (anonymous):

On the right side to \[(x-16)^{\frac{ 1 }{2 }}\]

OpenStudy (anonymous):

you have to change to say \(u=x-16\) to make it look like expanding \(\sqrt{16+u}\) at \(u=0\) that is where the expansion works then algebra to get \[\sqrt{16+u}=4\sqrt{1+\frac{u}{16}}\]

OpenStudy (anonymous):

My mind is blown. What I have thus far is incorrect? So where do I restart?

OpenStudy (anonymous):

now you can expand as before \[(16+z)^{1/2}=4\cdot\left(1+\frac u{16}\right)^{1/2}=\\ =4\cdot\sum \binom{1/2}n\frac{u^n}{16^n}\,.\]

OpenStudy (anonymous):

it is a pain to do it this way, might actually be easier to do it by hand

OpenStudy (anonymous):

but if you are going to use that formula for the binomial, you need it to look like \[(1+x)^{\frac{1}{2}}\]

OpenStudy (anonymous):

actually, since you probably want to leave it in sigma notation and not actually compute the coefficients you are done at this step

OpenStudy (anonymous):

or at this step \[4\cdot\sum \binom{1/2}n\frac{(x-16)^n}{16^n}\,.\]

OpenStudy (anonymous):

Once again where did the U Sub come from?

OpenStudy (anonymous):

Or just why did we introduce a U Sub?

OpenStudy (anonymous):

i see from your paper you were missing the 16 in the denominator using the this method the gimmick is to make it look like \((1+z)^{1/2}\) for some \(z\)

OpenStudy (anonymous):

that is the form it has to be in to use the binomial expansion

OpenStudy (anonymous):

Awesome thanks!

OpenStudy (anonymous):

yw, i think all steps are there